OpenStudy (anonymous):

Factor completely: 49x^2-25

OpenStudy (anonymous):

(7x-5)(7x+5)

OpenStudy (anonymous):

Difference of squares: 49x^2 - 25 = (7x +5) (7x-5)

OpenStudy (anonymous):

It is the difference of two squares. 49x^2 = (7x)^2 25 = 5^2 It can therefore be factorised as: (7x - 5)(7x + 5).